OpenStudy (anonymous):

(HELP PLEASE) (I WILL GIVE A MEDAL) Create your own word problem that can be solved using a system of equations. You will need to type the problem, the system, and the solution to receive full credit.

OpenStudy (anonymous):

I got out of surgery less than a week ago and I haven't been holding up well. Soon I will have to be hospitalized. I just need help on this. Please.

OpenStudy (texaschic101):

I do not know what problems were used in your notes or activities

OpenStudy (anonymous):

oh. it doesn't matter. sorry

OpenStudy (tkhunny):

We sold two kinds of tickets, Cheap Seats and Horrible Seats. Total tickets sold was 864 Total income from ticket sales was $4,521.22 What were the two ticket prices? It will not come out pretty, but there is one.

OpenStudy (texaschic101):

I wonder how many people bought the horrible seats....lol

OpenStudy (tkhunny):

Sometimes, it's all one can afford. Sad, but true.

OpenStudy (anonymous):

O.O can you give me the answers and stuff pease. i have to have work woth it.

OpenStudy (anonymous):

*with

OpenStudy (tkhunny):

Give it a go! Let's see what you can manage.

OpenStudy (anonymous):

I've tried..

undeadknight26 (undeadknight26):

I did this not too long ago...but im really tired right now...I really hope you will get better! The world will be chaos for guys like me without cute gals like you ;-; *thumbs up* Hope i at least got a smile!

OpenStudy (anonymous):

I don't understand any of this..

undeadknight26 (undeadknight26):

Wait...this is 9th right?

OpenStudy (anonymous):

I think so..

undeadknight26 (undeadknight26):

Well james tom kanta saizo sunabozu black tomahawk terry jerry bob fredrick ricky mason kelly and black is my name! I got you! let me just try to remember what a system of equations is!

OpenStudy (anonymous):

o.o.. long name? and thank you

undeadknight26 (undeadknight26):

Please call me...tom...or bob the turtle...or simply Fredrick.

undeadknight26 (undeadknight26):

What do they mean by system. Would they like a picture of the system?

OpenStudy (anonymous):

I will do my best to remember Tom.

OpenStudy (anonymous):

I need to type up an example and and it has to be solvable

undeadknight26 (undeadknight26):

Please give me 4 of more of your favorite numbers...lets try to keep the numbers below...10!

OpenStudy (anonymous):

I can do that! hmm.. 2 5 8 9

undeadknight26 (undeadknight26):

Let me slightly modify them for an easy solution! y = 6x - 8 y = 2 -4x

undeadknight26 (undeadknight26):

The solution is 1, -2

undeadknight26 (undeadknight26):

This is easily checked by: -2 = 6(1) - 8 -2 = 2 - 4(1)

OpenStudy (anonymous):

How would I put that in a word problem though?

undeadknight26 (undeadknight26):

*itches head* does i has to come up with every thing? XD lol im jk...Hmm...Lets say... Bob the turtle wanted to find out which part of water was the coolest. His friend Fredrick the Pheasant flew up and gave him the formulas y = 6x - 8, y = 2 -4x to find the coolest part of the water. How would bob the turtle solve this?

undeadknight26 (undeadknight26):

Hope you get better! If you need anything please do tag/pm me. Although i may be gaming. Sayonara!

OpenStudy (anonymous):

o.o... thank you haha

undeadknight26 (undeadknight26):

Sowwy i crazy.

OpenStudy (anonymous):

It's cool. I like the word problem. It's creative

undeadknight26 (undeadknight26):

bob the turtle is real. Fredrick was a frog...But pheasants are more graceful.

OpenStudy (anonymous):

haha awe. thats so cute!

undeadknight26 (undeadknight26):

Have you ever seen a pheasant? i have always wanted to see one...bob the turtle is a random turtle me and my friends rescued from a random Chinese restaurant and Fredrick teh frog followed me and my friends around for 4 straight days... Weird right?

OpenStudy (anonymous):

haha yeah and yes. they have cute beaks